I have a problem solving this word problem Thank you

A plane took 1 hr longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?

Answer by checkley75(3666)

560/(X+1)=480/X NOW CROSS MULTIPLY
560X=480X+480
560X-480X=480
80X=480
X=480/80
X=6 HOUR IS THE FLYING TIME FOR THE 480 MPH PLANE.
480/6=80 MPH.